OpenStudy (anonymous):

Verify the identity. cos (x + pi/2) = -sin x

OpenStudy (anonymous):

@gorv

OpenStudy (gorv):

pi/2=90 degree

OpenStudy (gorv):

now if we add something it will go in second quaderent

OpenStudy (gorv):

in second quad cos give negative value

OpenStudy (gorv):

cos(pi/2+x)=-sinx

OpenStudy (anonymous):

\[\cos \left( A+B \right)=\cos A \cos B-\sin A \sin B\] \[\cos \frac{ \pi }{ 2 }=0,\sin \frac{ \pi }{ 2 }=1\]

OpenStudy (gorv):

or u can use formula cos(a+b)

OpenStudy (anonymous):

so I'm having trouble figuring out where to plug things into but i'll try cos A + B = cos pi/2 cos 0 - sin pi/2 sin 1 I don't think any of that was right @gorv